OpenStudy (anonymous):

2. Write as an exponent: Log2 ½ = -1

OpenStudy (zarkon):

\[\Large \log_{a}(b)=c\] \[\Large a^c=b\]

OpenStudy (anonymous):

2^-1 = 1/2

OpenStudy (anonymous):

@pooja195

jhonyy9 (jhonyy9):

so you know that 1/2 = 2^(-1) what mean that you can rewriting your exercise log2 1/2 = log2 2^(-1) = (-1)log2 2 = -1 because you know that log2 2 = 1 hope this will help you understanding sure easy